# AVG-LLaVA: A Multimodal Large Model with Adaptive Visual Granularity

### v1

@@ Abstract
Recently, when dealing with high-resolution images, dominant LMMs usually divide them into multiple local images and one global image, which will lead to a large number of visual tokens. In this work, we introduce AVG-LLaVA, an LMM that can adaptively select the appropriate visual granularity based on the input image and instruction. This approach not only reduces the number of visual tokens and speeds up inference, but also improves the overall model performance. Specifically, we introduce the following modules based on LLaVA-NeXT: (a) a visual granularity scaler that includes multiple pooling layers to obtain visual tokens with different granularities; (b) a visual granularity router, which includes a Transformer layer, an MLP layer, and a voter layer, used to select the appropriate visual granularity based on the image and instruction. Furthermore, we propose RGLF, a novel training paradigm that aims at aligning the granularity predicted by the router with the preferences of the LMM, without the need for additional manually annotated data. Extensive experiments and analysis show that AVG-LLaVA achieves superior performance across 11 benchmarks, as well as significantly reduces the number of visual tokens and speeds up inference (e.g., an 85.3% reduction in visual tokens and a 2.53× increase in inference speed on the AI2D benchmark). The code will be available at https://github.com/DeepLearnXMU/AVG-LLaVA.
@@ 1 Introduction
Recently, the field of artificial intelligence (AI) has witnessed a significant advancement in Large Multimodal Models (LMMs) (OpenAI, 2023b; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023a; 2024a), marking a further step toward artificial general intelligence (AGI). Most existing LMMs follow the structure of LLaVA (Liu et al., 2023a; 2024a), which includes a vision encoder to embed images into visual tokens and a connector to map them into the LLM’s word embedding space. Subsequently, these visual tokens are fed into a Large Language Model (LLM) (Touvron et al., 2023; OpenAI, 2023a; Chiang et al., 2023) for multimodal understanding and reasoning, alongside the word embeddings.
@@ 2 Related Work
Large Multimodal Models.
@@
High-Resolution LMMs.
@@
Vision Token Reduction for LMMs.
@@ 3.1 Model Architecture
High-Resolution Image Encoding.
@@
Visual Granularity Scaler.
@@
Visual Granularity Router.
@@ 3.2 Multi-stage Training
Stage 2: Visual Instruction Tuning.
@@
Stage 3: Multi-Granularity Visual Instruction Tuning.
@@
Stage 4: Ranking Granularity to Align LMM Feedback.
@@ 4.1 Settings
Training Datasets.
Note that in this work, we mainly focus on investigating the effectiveness of automatic visual granularity selection in reducing the number of visual tokens and improving model performance. Therefore, during the first stage, we also use CC-595K image-text pairs for model training, as implemented in LLaVA-NeXT (Liu et al., 2024b). In the subsequent training stages, we also hope to use the same data as LLaVA-NeXT. However, the real user interaction data used for visual instruction fine-tuning in LLaVA-NeXT are not open-sourced, so we opt to extract 200K samples from ALLaVA (Chen et al., 2024a) dataset as a substitute. Although LLaVA-NeXT replaces TextVQA (Singh et al., 2019) with DocVQA (Mathew et al., 2021) and SynDog-EN (Kim et al., 2022), the TextVQA has already been included in the training data of most existing LMMs. Consequently, we choose to retain it to ensure a fair comparison with other models [^2] Our data recipe refers to https://github.com/xiaoachen98/Open-LLaVA-NeXT.. In total, the visual instruction fine-tuning data we use contains 1M image-text pairs.
Implementation Details.
@@ 4.2 Main Results
General Multimodal Benchmarks.
@@ 4.5 Ablation Study
Adaptive Visual Granularity vs. Fixed Visual Granularity.
@@
Router Granularity Selection vs. Random Granularity Selection.
@@
Impact of Router Input.
@@
Impact of Granularity Range.
@@
Impact of Router Training Methods.
@@
Importance of Ranking Granularity.

### v2

# AVG-LLaVA: A Large Multimodal Model with Adaptive Visual Granularity

@@ Abstract
Recently, when dealing with high-resolution images, dominant large multimodal models (LMMs) usually divide them into multiple local images and one global image, which will lead to a large number of visual tokens. In this work, we introduce AVG-LLaVA, an LMM that can adaptively select the appropriate visual granularity based on the input image and instruction. This approach not only reduces the number of visual tokens and speeds up inference, but also improves the overall model performance. Specifically, we introduce the following modules based on LLaVA-NeXT: (a) a visual granularity scaler that includes multiple pooling layers to obtain visual tokens with different granularities; (b) a visual granularity router, which includes a Transformer layer, an MLP layer, and a voter layer, used to select the appropriate visual granularity based on the image and instruction. Furthermore, we propose RGLF, a novel training paradigm that aims at aligning the granularity predicted by the router with the preferences of the LMM, without the need for additional manually annotated data. Extensive experiments and analysis show that AVG-LLaVA achieves superior performance across 11 benchmarks, as well as significantly reduces the number of visual tokens and speeds up inference (e.g., an 85.3% reduction in visual tokens and a 2.53$\times$ increase in inference speed on the AI2D benchmark). The code will be available at https://github.com/DeepLearnXMU/AVG-LLaVA.
@@ 1 Introduction
Recently, the field of artificial intelligence (AI) has witnessed a significant advancement in large multimodal models (LMMs) (OpenAI, 2023b; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023a; 2024a), marking a further step toward artificial general intelligence (AGI). Most existing LMMs follow the structure of LLaVA (Liu et al., 2023a; 2024a), which includes a vision encoder to embed images into visual tokens and a connector to map them into the word embedding space. Subsequently, these visual tokens are fed into a large language model (LLM) (Touvron et al., 2023; OpenAI, 2023a; Chiang et al., 2023) for multimodal understanding and reasoning, alongside the word embeddings.
@@ 2 Related Work
Large Multimodal Models
@@
High-Resolution LMMs
@@
Vision Token Reduction for LMMs
@@ 3.1 Model Architecture
High-Resolution Image Encoding
@@
Visual Granularity Scaler
@@
Visual Granularity Router
@@ 3.2 Multi-stage Training
Stage 2: Visual Instruction Tuning
@@
Stage 3: Multi-Granularity Visual Instruction Tuning
@@
Stage 4: Ranking Granularity to Align LMM Feedback
@@ 4.1 Settings
Training Datasets
Note that in this work, we mainly focus on investigating the effectiveness of automatic visual granularity selection in reducing the number of visual tokens and improving model performance. Therefore, during the first stage, we also use CC-595K image-text pairs for model training, as implemented in LLaVA-NeXT (Liu et al., 2024b). In the subsequent training stages, we also hope to use the same data as LLaVA-NeXT. However, the real user interaction data used for visual instruction fine-tuning in LLaVA-NeXT are not open-sourced, so we opt to extract 200K samples from ALLaVA (Chen et al., 2024a) dataset as a substitute. Although LLaVA-NeXT replaces TextVQA (Singh et al., 2019) with DocVQA (Mathew et al., 2021) and SynDog-EN (Kim et al., 2022), the TextVQA has already been included in the training data of most existing LMMs. Consequently, we choose to retain it to ensure a fair comparison with other models. [^2] Our data recipe refers to Open-LLaVA-NeXT (Lin & Long, 2024) In total, the visual instruction fine-tuning data we use contains 1M image-text pairs.
Implementation Details
@@ 4.2 Main Results
General VQA Benchmarks
Text-oriented VQA Benchmarks
General Multimodal Benchmarks
@@ 4.5 Ablation Study
Adaptive Visual Granularity vs. Fixed Visual Granularity
@@
Router Granularity Selection vs. Random Granularity Selection
@@
Impact of Router Input
@@
Impact of Granularity Range
@@
Impact of Router Training Methods
@@
Importance of Ranking Granularity